solve it with step by step please and how each number is calculated

Southwestern University (SWU), a large state college in Stephenville, Texas enrolls close to 20,000 students. It is the dominant force in the small city, with more students during fall and spring than permanent residents.

A longtime football powerhouse, the University s a member of the Big Eleven conference and is usually in the top 20 in college football rankings. To bolster its chances of reaching the elusive and long-desired number-one ranking, in 2013 the University hired the legendary Billy Bob Dillon as its head coach. Although the number-one ranking remained out of reach, attendance at the six Saturday home games each year increased. Prior to Dillon’s arrival, attendance generally averaged 30,000–35,000. Season ticket sales bumped up by 15,000 just with the announcement of the new coach’s arrival. Stephenville and the University were ready to move to the big time!

With the growth in attendance came more fame, the need for a bigger stadium, and more complaints about seating, parking, long lines, and concession stand prices. The University’s president was concerned not only about the cost of expanding the existing stadium versus building a new stadium but also about the ancillary activities. He wanted to be sure that these various support activities generated revenue adequate to pay for themselves. Consequently, he wanted the parking lots, game programs, and food service to all be handled as profit centers. At a recent meeting discussing the new stadium, the University’s president told the stadium manager to develop a break-even chart and related data for each of the centers. He instructed the stadium manager to have the food service area break-even report ready for the next meeting. After discussion with other facility managers and his subordinates, the stadium manager developed the following table showing the suggested selling prices, his estimate of variable costs, and his estimate of the percentage of the total revenues that would be expected for each of the items based on historical sales data.

Item

Selling Price/Unit

$

Variable Cost/Unit

$

Percent Revenue

%

Soft drink

1.00

0.45

15%

Coffee

3.00

0.60

30%

Hot dogs

3.50

0.85

25%

Burgers

4.50

1.50

18%

Misc. Snacks

2.00

0.90

12%

The stadium manager’s fixed costs are interesting. He estimated that the prorated portion of the stadium cost would be as follows: salaries for food services at $360,000 ($60,000 for each of the six home games); 2,400 square feet of stadium space at $ 8 per square foot per game; and six people per booth in each of the six booths for 5 hours at $15 an hour. These fixed costs will be proportionately allocated to each of the products based on the percentages provided in the table. For example, the revenue from soft drinks would be expected to cover 20 % of the total fixed costs.

The stadium manager wants to be sure that he has a number of things for the University President: (1) the total fixed cost that must be covered at each of the games; (2) the portion of the fixed cost al- located to each of the items; (3) what his unit sales would be at breakeven for each item—that is, what sales of soft drinks, coffee, hot dogs, hamburgers, and snacks are necessary to cover the portion of the fixed cost allocated to each of these items; (4) what the dollar sales for each of these would be at these break-even points; and (5) realistic sales estimates per attendee for attendance of 60,000 and 35,000. (In other words, he wants to know how many dollars each attendee is spending on food at his projected break-even sales at present and if attendance grows to 60,000.) He felt this last piece of information would be helpful to understand how realistic the assumptions of his model are, and this information could be compared with similar figures from previous seasons.

Calculate                                                                                                                                  

  1. The total fixed cost per game and the allocated cost to each food item              
  2. Calculate the Break-Even-Point (BEP) for each of these items                                 
  3. Determine the total sales for each item that is required to each BEP                    
  4. Write your conclusion whether we break even or not if the attendance               

is 35000 people or 60000 people.

  1. Prepare a brief report for the University President that covers the items noted.
Answer #1

Solution to part a- Total fixed cost per game and the allocated cost to each food item:

Computation of total fixed cost per games-

Salaries for food services- 60,000      Stadium rent (2400sqft*$8)-19,200 Booth Charges (6 people*6 booth*5 hours*$15)-2,700

Total fixed cost:$81,900/game

Computation of allocated cost to each food item-

Soft Drink-81900*15% =12285 Coffee-81900*30%=24570 Hot dogs-81900*25%=20475 Burgers-81900*18%=14742 Misc. Snacks-81900*12%=9828

Solution to part b.

Computation of BEP for each food item-

Item Sale price(a) Variable Cost(b) Contribution(a-b) Fixed Cost(c) Break Even Point(c/b)
Soft Drink 1.00 .45 .55 12285 22337
Coffee 3.00 .60 2.40 24570 10238
Hot Dogs 3.50 .85 2.65 20475 7727
Burger 4.50 1.50 3.00 14742 4914
Misc 2.00 .90 1.10 9828 8935

Note:All calulations of BEP have been rounded up

Solution to part -c : Total sales of each item to acheive BEP sales-

Item Sale price(a) Break Even Point(b) Break Ecen Sales (a*b)
Soft Drink 1.00 22337 22337
Coffee 3.00 10238 30714
Hot Dogs 3.50 7727 27044.50
Burger 4.50 4914 22113
Misc 2.00 8935 17870

Solution to part d-Break Even if attendance is 35000 people or 60000 people

Computation of overall break even point-

Expected contribuiton should be equal to fixed costs =81900

Weighted Averege Contribution:

Items Contribution(a) Sales Mix%(b) Total Contribution(c:a*b)
Soft Drink .55 15 .0825
Coffee 2.40 30 .72
Hot Dog 2.65 25 .106
Burger 3 18 .54
Misc 1.1 12 .132

Weighted Average Contribution=Total of c=1.508

Overall BES=81900/1.508=54311attendees

Therefore at current level of 35000 people the university is not breaking even as the number of attendees is lower than 54311,however at 60000 people the university shall be able to break even.

Solution to part e-

Break Even Report

At the current level of activity the fixed costs consisting of food staff salaries booth charges and stadium rent amount to 81900 per game and the overall break even is reached when the number of attendees reach 54311 and profits starts when attendees are more than 54311.Now, at each level of food items,after allocating the fixed costs,the break even points are 22337 units for soft drinks,10238 for coffee,7727 for hot dogs,4914 for burger and 8935 for misc snacks.